hi, would you please help me solve this multiple choice question?:
one leg of a right triangle is twice as long as the other leg. the area of the triangle is 36 square feet. what is the length of the shorter leg? the choices are:
a) 5 ft b) 6 ft c) 7 ft d) 8 ft
i know the answer is 6 ft but i'd like to know how you get this answer. thank you!

Answer by amarjeeth123(569)

Let the length of the shorter leg be x.
Then the length of the longer leg is 2x.
Area of the triangle=36 square feet
1/2*base*height=36
1/2*x*2x=36
x^2=36
x=6
The solution is x=6 feet.